# Nonequilibrium steady states in multi-bath quantum collision models

### v1

@@ Abstract
Collision models provide a simple and versatile setting to capture the dynamics of open quantum systems. The standard approach to thermalisition in this setting involves an environment of independent and identically-prepared thermal qubits, interacting sequentially for a finite duration $\Delta t$ with the system. We compare this to a two-bath scenario in which collisional qubits are prepared in either their ground or excited states and the environment temperature is encoded in system-environment couplings. The system reaches the same thermal steady state for both settings as $\Delta t\to 0$, although even in this limit they describe fundamentally different physical processes, with the two-bath setup yielding a nonequilibrium state with finite heat currents. Non-Markovian dynamics arise when intra-environment interactions in either setting are introduced. Here, the system in the single-bath setup again reaches a steady state at the canonical temperature of the bath, but the nonequilibrium steady state of the two-bath setup tends to a different temperature due to the generation of strong system-environment and intra-environment correlations. The two-bath setting is particularly suited to studying quantum trajectories, which are well-defined also for the non-Markovian case. We showcase this with a trajectory analysis of the heat currents within a two-point measurement scheme. Our results provide insights into the dynamics and thermodynamics of thermalisation towards nonequilibrium steady states and the role of non-Markovian interactions.
@@ I Introduction
A framework for understanding how a quantum system evolves when in contact with an environment is critical; both in terms of the practical development of quantum technologies and in addressing long-standing open questions regarding the dynamics of complex systems. Regarding the former noise characterisation in prototype quantum devices is crucial to overcome scalability issues Filippov et al. (2024), while the latter touches on major conceptual gaps in our understanding, e.g. determining the quantum-to-classical transition Zurek (2003) or how complex systems thermalise Deutsch (2018). It is not surprising then that a range of tools have been developed to simulate open quantum systems Purkayastha et al. (2021); Ferracin et al. (2024); Landi et al. (2022); Breuer and Petruccione (2002); Brenes et al. (2020). Within this plethora of techniques, collision models Ciccarello et al. (2022); Campbell and Vacchini (2021); Rau (1963); Scarani et al. (2002); Ziman et al. (2002) or repeated interaction schemes Strasberg et al. (2017), stand as a particularly versatile approach having been used to reconcile apparent thermodynamic inconsistencies present in other approaches De Chiara et al. (2018), explore how classicality emerges from an otherwise quantum dynamics Campbell et al. (2019); Ryan et al. (2022), and efficiently cool quantum systems Taranto et al. (2020); Langbehn et al. (2025).
@@
The simplest formulation of a collision model involves imagining the environment as an infinite collection of identically prepared constituents, here after referred to as “auxiliary units” and labelled $A$, with which the system, $S$, sequentially interacts. Depending on the precise form of the $S$-$A$ interaction, and the initial states of the auxiliary units, this model accurately captures a range of Markovian dynamics. In particular, for so-called “thermal operations” it leads to thermalisation Scarani et al. (2002), while the model also more generally leads to homogenisation Ziman et al. (2002). However, varying the $S$-$A$ interaction can lead to a range of non-equilibrium steady states (NESS), whose characterisation has recently become the focus of several works Guarnieri et al. (2020); Heineken et al. (2021); Tian et al. (2021); Prositto et al. (2025); Palafox et al. (2022); Santiago-García et al. (2025). While assuming that all the auxiliary units are identically prepared is well justified, particularly in order to simplify simulation, this is clearly not a very realistic model. For instance, a more precise microscopic description of a thermal environment would assume that the system interacts with auxiliary units whose initial state is drawn in accordance with a distribution of energy that correspond to a given temperature. Several works have explored relaxing the assumption of identical auxiliary units and deterministic collisional events, leveraging scattering techniques to understand a wide class of open system dynamics Ehrich et al. (2020); Jacob et al. (2021); Tabanera-Bravo et al. (2023). The flexibility of collision models is clearly one of their inherent strengths. From the preceding discussion it is clear that collision models provide a useful approach to simulate how a system reaches the same, possibly thermal or equilibrium, steady state from different microscopic models.
@@
This serves as the starting point for our analysis. We model a simple thermalisation process via collision models for two settings: the ‘typical’ situation where the auxiliary units are assumed to be all identical thermal states and a second setting where the system interacts simultaneously with two collisional baths with different initial energies, and these energies dictate the strength of the $S$-$A$ interaction. While these approaches are known to precisely recover the expected Markovian dynamics as described by the standard GKSL master equation, we show that they correspond to physically distinct dynamics, with the latter setting clearly driving the system to a non-equilibrium steady state. Through a simple modification of the basic collision model, we examine how the introduction of non-Markovianity to the dynamics significantly changes the steady state properties observed, with the establishment of strong system bath correlations playing a particularly important role. Finally, we explore the impact of introducing projective measurements on the auxiliary units using a two-point measurement scheme to recover the average heat current in the presence of non-Markovianity.
@@ II Markovian Quantum Collision models
II.1 Single-bath homogenizing model (setting I)
@@
Collision models are a flexible framework for studying open system dynamics Ciccarello et al. (2022); Campbell and Vacchini (2021); indeed they can be used to simulate any time-continuous open system dynamics described by a Markovian quantum master equation Ziman et al. (2005). A simple setting of wide physical relevance is a two-level system with Hamiltonian, $H_{S}=\tfrac{\omega_{S}}{2}\sigma_{z}$, thermalising with a bosonic environment described by the master equation Landi et al. (2024),
@@
where $\Gamma$ is the coupling strength and $\bar{N}\!=\!(e^{\beta\omega_{S}}-1)^{-1}$ is the Bose-Einstein distribution for a bath at (inverse) temperature $\beta\!=\!1/T$, as depicted in Fig. 1(a). The two dissipators $\mathcal{D}[\sigma^{-}],\mathcal{D}[\sigma^{+}]$, with $\sigma^{-},\sigma^{+}$ the Pauli lowering and raising operators, represent the emission and absorption of thermal photons, respectively. The steady state of Eq. (4) is given by the Gibbs state,
@@
The dynamics described by Eq. (4) can be stroboscopically simulated using a collision model Rau (1963); Scarani et al. (2002); Ciccarello et al. (2022), wherein the environment is viewed as an infinite collection of auxiliary units, each modeled as a qubit, identically prepared in a thermal state, i.e. $\eta_{n}=\rho^{\text{th}}$ given by Eq. (5), with the free evolution of each unit given by $H_{A_{n}}=\tfrac{\omega_{A_{n}}}{2}\sigma^{z}$. For simplicity, in the rest of this work we shall always assume the system and auxiliary units are on resonance, i.e. $\omega_{S}\!=\!\omega_{A_{n}}\!\equiv\!\omega$. This setting is illustrated in Fig. 1(b). In order to match the dynamics given by Eq. (4), we must fix the interaction Hamiltonian between the $S$ and $A_{n}$ to be,
@@
with $J=\sqrt{{\frac{\Gamma(2\bar{N}+1)}{\Delta t}}}$. Scaling the interaction strength by $1/\sqrt{\Delta t}$ is necessary to ensure that Eq. (4) is recovered in the limit $\Delta t\to 0$. Throughout this work we will refer to this as ‘setting I’. Henceforth, we shall move to the interaction picture and neglect the free evolution of the system and auxiliary units. This can be done due to the fact that sum of the free Hamiltonians commute with the interaction Hamiltonian, i.e. $\left[H_{S}+H_{A_{n}},H_{I}\right]=0$  Lorenzo et al. (2017). This fact also means there is no switching work required to stoboscopically turn on and off the system-bath interaction De Chiara et al. (2018); Guarnieri et al. (2020). Under these conditions, the unitary in Eq. (1) corresponds to a partial-SWAP operation Scarani et al. (2002); McCloskey and Paternostro (2014); Campbell et al. (2018),
where $\mathcal{W}$ is the SWAP operator,
$\mathcal{W}=\begin{pmatrix}1&0&0&0\\ 0&0&1&0\\ 0&1&0&0\\ 0&0&0&1\\ \end{pmatrix}.$ (8)
Figure 1: Schematic depiction of the approaches to model open quantum system dynamics considered in this work. (a) Typical open system dynamics where the system is interacts with a bath, generally assumed to be much larger than the system. Under certain assumptions this can be modeled using master equations. (b) Setting I: Collision models involve partitioning the environment into smaller constituent parts all prepared in the same initial state, $\eta_{i}$. The system interacts with a given environmental unit via some interaction with strength $J$. For interactions of the form Eq. (6) this collision model leads to homogenisation and the system reaches the steady state $\rho_{\infty}\!=\!\eta_{i}$. If $\eta_{i}$ are thermal states, this dynamics can be shown to be equivalent to the master equation, Eq. (4). (c) Setting II: Two bath collision model. The system interacts simulataneously with two independent collisional baths, where the constituents are initially in pure states |$\ket{0}$ and |$\ket{1}$, respectively, whose interaction strengths encode the effective temperature of the bath. In the limit of small interaction times, the system reaches the same steady state as setting I.
@@
This set-up is a standard example of quantum homogenization, wherein the steady state of $S$ converges to the initial auxiliary state, i.e. $\rho_{\infty}\!=\!\eta_{i}$, for any choice of initial state for the auxiliary units $A_{n}$ Ziman et al. (2002). When $S$ sequentially interacts with $A_{n}$ via partial swap operations, its dynamics are described by a contractive completely positive map, guaranteeing convergence toward $\eta$ as the unique fixed point. This homogenization process is thermodynamically consistent and captures key features of thermalization. Importantly, in the limit of infinitely many collisions, the system not only reaches equilibrium but also asymptotically factorizes from the environment, resulting in a product state without residual system-environment correlations Cusumano et al. (2018).
@@ II.2 Two-bath model (setting II)
An alternative approach to simulating the time-continuous dynamics of Eq. (4) involves allowing $S$ to interact with two collisional baths, one in which all the auxiliary units are in the |$\ket{0}$ state and one in which they are all in the |$\ket{1}$ state, where |$\ket{0}$ and |$\ket{1}$ corresponds to the ground and excited states of $A_{n}$, respectively. This setting is illustrated in Fig. 1(c). Throughout this work we will refer to this as ‘setting II’. The system interacts with a qubit from each bath simultaneously via an interaction Hamiltonian $H_{I}^{(i)}$ during a given collision, where we have introduced the superscript $i=0,1$ to label interactions with the baths |$\ket{0}$ and |$\ket{1}$, respectively. The interaction Hamiltonians take the same form as Eq. (6) but with different interaction strengths, $J^{(i)}$, which are given by
@@
In contrast to setting I, in this case the system will be driven to a non-equilibrium steady state (NESS), due to the simultaneous interaction with multiple inequivalent baths. This is because, irrespective of whether the system thermalizes or not, there will always be a constant flow of energy from the |$\ket{1}$ bath to the |$\ket{0}$ bath.
The unitary in Eq. (1) is also no longer given simply by a partial swap, and thus the system cannot homogenize. Similar to setting I however, we have $\left[H_{S}+H_{A_{n}},H_{I}^{(0)}+H_{I}^{(1)}\right]=0$ and so the free evolutions of the system and auxiliary units can again be neglected.
@@ II.3 Steady States in the Markovian Limit
We start our analysis by noting that settings I and II both simulate the dynamics of a system in contact with a thermal bath. Indeed, the two settings are equivalent in the limit of $\Delta t\rightarrow 0$, where they both recover the dynamics and steady state that would be expected under the time continuous evolution given by Eq. (4). Furthermore, it is evident that they are equivalent in the zero temperature limit, $\beta\rightarrow\infty$. This is an obvious consequence of the fact that, in this limit, $J^{(1)}\!\to\!0$ and thus the two bath case reduces to the zero temperature limit of setting I. However, as alluded to previously, and as we will show explicitly below, setting II generally drives the system to a NESS.
The analytical form of the steady state in setting II is readily accessible, although too cumbersome to provide any direct insight. Nevertheless, it is diagonal in the system’s energy eigenbasis and, since we are considering a simple two-level system, it can therefore be assigned an effective temperature, $\beta_{e}$. This is formally defined as follows: as the state is diagonal, it can always be written in the form (5), with an effective parameter $g_{e}$. The effective temperature is then defined as
$\beta_{e}=\frac{1}{\omega}\log\left(\frac{1+g_{e}}{1-g_{e}}\right).$ (11)
As stated, the explicit formula of $g_{e}$ or $\beta_{e}$ is too cumbersome to write down. But, to illustrate the idea, in the case of very small $\Delta t$ we have
It therefore deviates from the equilibrium value $g=(2\bar{N}+1)^{-1}$ by a term that is proportional to $\Delta t$.
In Fig. 2 we show the deviation from the “canonical” steady state temperature, $\beta$, by calculating $\Delta\beta=\beta_{e}-\beta$. We clearly observe that in the limit of short collisions ($\Delta t\to 0$), the system is driven to the expected steady state, i.e. $\beta_{e}\!=\!\beta$, as predicted by Eq. (12). Increasing $\Delta t$, corresponding to longer interaction times during the collisions, leads to a substantial increase in $\Delta\beta$, with this deviation becoming more pronounced for increasing temperatures. Larger $\beta$ leads to larger deviations and we consistently find that $\Delta\beta\!>\!0$, indicating that the effective temperature that the system is driven to is cooler than expected. Considering each bath individually, it is clear that the $A_{n}^{(0)}$’s initialised in |$\ket{0}$ are at zero temperature, while the $A_{n}^{(1)}$’s in the |$\ket{1}$ can be considered as systems at negative infinite temperature. As the system is simulataneously in contact with both baths it follows that it must reach a NESS.
Maintaining the system in such a NESS requires heat currents to be flowing. Because the unitary is strictly energy preserving, there is no work associated with switching on and off the collisional unitaries De Chiara et al. (2018); Guarnieri et al. (2020). Hence all energy changes can be considered as heat exchanges. Focusing on the zero temperature bath in setting II, within the collision model setting we can readily calculate this via
@@
where $\tilde{\rho}_{A_{n}^{(0)}}$ is the state of the $A_{n}^{(0)}$ auxiliary unit after the system-environment collision has taken place. A similar expression holds for the $A_{n}^{(1)}$’s and, in fact, we find is exactly equal and opposite to Eq. (13), which is a direct consequence of strict energy conservation. The inset of Fig. 2 shows the steady state heat flux into the environment, i.e. $\frac{\Delta\mathcal{Q}}{\Delta t}\!\equiv\!\frac{\Delta\mathcal{Q}^{(0)}(n)}{\Delta t}$ for $n\!\rightarrow\!\infty$. We see that it is non-zero for all values of $\Delta t$. Smaller values of $\beta$ require larger heat fluxes to maintain the NESS. This is in line with expectation since in this regime the system is being driven to a state with a high effective temperature. It is therefore far from equilibrium from both of the respective collisional baths leading to substantial energy exchanges. Larger $\beta$ leads to a reduced heat flux since this regime tends to the zero temperature single bath situation captured by setting I. It is worth noting that the heat flux vanishes for all choices of $\beta$ in setting I Guarnieri et al. (2020).
Figure 2: Steady states in the Markovian limit. Main panel: Deviation of the effective temperature from the canonical temperature, $\Delta\beta$, with increasing collision time-step $\Delta t$ for different canonical temperatures $\beta\!=\!2$ (dashed, blue; lower temperature) and $\beta\!=\!0.5$ (solid, red; higher temperature). Inset: Non-equilibrium steady state heat current, $\Delta\mathcal{Q}/\Delta t$ for the initially low energy (cold) bath, $A_{n}^{(0)}$.

### v2

@@ Abstract
Collision models provide a simple and versatile setting to capture the dynamics of open quantum systems. The standard approach to thermalisition in this setting involves an environment of independent and identically-prepared thermal qubits, interacting sequentially for a finite duration $\Delta t$ with the system. We compare this to a two-bath scenario in which collisional qubits are prepared in either their ground or excited states and the environment temperature is encoded in system-environment couplings. The system reaches the same thermal steady state for both settings, although even in this limit they describe fundamentally different physical processes, with the two-bath setup yielding a nonequilibrium state with finite heat currents. Non-Markovian dynamics arise when intra-environment interactions in either setting are introduced. Here, the system in the single-bath setup again reaches a steady state at the canonical temperature of the bath, but the nonequilibrium steady state of the two-bath setup tends to a different temperature due to the generation of strong system-environment and intra-environment correlations. The two-bath setting is particularly suited to studying quantum trajectories, which are well-defined also for the non-Markovian case. We showcase this with a trajectory analysis of the heat currents within a two-point measurement scheme. Finally, we consider how our results are impacted when the system-environment interaction leads to strict homogenisation. Our results provide insights into the dynamics and thermodynamics of thermalisation towards nonequilibrium steady states and the role of non-Markovian interactions.
@@ I Introduction
A framework for understanding how a quantum system evolves when in contact with an environment is critical; both in terms of the practical development of quantum technologies and in addressing long-standing open questions regarding the dynamics of complex systems. Regarding the former noise characterisation in prototype quantum devices is crucial to overcome scalability issues ManiscalcoScalability, while the latter touches on major conceptual gaps in our understanding, e.g. determining the quantum-to-classical transition ZurekRMP or how complex systems thermalise Deutsch2018. It is not surprising then that a range of tools have been developed to simulate open quantum systems PReB; Ferracin2024; LandiRMP; Breuer2002; GooldPRX. Within this plethora of techniques, collision models CICCARELLO20221; Campbell2021EPL; cusumano2022quantum; Rau1963; ScaraniPRL2002; Homogenisation or repeated interaction schemes StrasbergPRX, stand as a particularly versatile approach having been used to reconcile apparent thermodynamic inconsistencies present in other approaches DeChiaraNJP, explore how classicality emerges from an otherwise quantum dynamics CampbellPRA2019; Ryan2022, explore synchronization in quantum systems karpat2019pra, and efficiently cool quantum systems HuberCooling; KochCooling.
@@
The simplest formulation of a collision model involves imagining the environment as an infinite collection of identically prepared constituents, here after referred to as “auxiliary units” and labelled $A$, with which the system, $S$, sequentially interacts. Depending on the precise form of the $S$-AA interaction, and the initial states of the auxiliary units, this model accurately captures a range of Markovian dynamics. In particular, for so-called “thermal operations” it leads to thermalisation ScaraniPRL2002, while the model also more generally leads to homogenisation Homogenisation. However, varying the $S$-AA interaction can lead to a range of non-equilibrium steady states (NESS), whose characterisation has recently become the focus of several works Guarnieri2020PLA; Heineken_NESS_Entanglement; Tian_Heat_Transfer_2021; Segal2025QST; Baris1; Baris2; Hammam2024quantum; cusumano2024structured; hammam2021optimizing; corr2025continuous. While assuming that all the auxiliary units are identically prepared is well justified, particularly in order to simplify simulation, this is clearly not a very realistic model. For instance, a more precise microscopic description of a thermal environment would assume that the system interacts with auxiliary units whose initial state is drawn in accordance with a distribution of energy that correspond to a given temperature KochCooling. Several works have explored relaxing the assumption of identical auxiliary units and deterministic collisional events, e.g. by leveraging scattering techniques Scattering1; Scattering2; Scattering3. The flexibility of collision models is clearly one of their inherent strengths that allows us to explore a range of other open system dynamics Lorenzo2015heat; cusumano2017interferometric; cusumano2018interferometric; pedram2022environment; Erbanni2023simulating; Garg2025simulating. From the preceding discussion it is clear that collision models provide a useful approach to simulate how a system reaches the same, possibly thermal or equilibrium, steady state from different microscopic models.
@@
This serves as the starting point for our analysis. We model a simple thermalisation process via collision models for two settings: the ‘typical’ situation where the auxiliary units are assumed to be all identical thermal states and a second setting where the system interacts simultaneously with two collisional baths with different initial energies, and these energies dictate the strength of the $S$-AA interaction. While these approaches are known to recover the expected Markovian dynamics as described by the standard GKSL master equation CICCARELLO20221; cusumano2022quantum (see the Appendix), we show that they correspond to physically distinct processes, with the latter setting clearly driving the system to a NESS. Through a simple modification of the basic collision model, we examine how the introduction of non-Markovianity to the dynamics significantly changes the steady state properties observed, with the establishment of strong system bath correlations playing a particularly important role. We explore the impact of introducing projective measurements on the auxiliary units using a two-point measurement scheme to recover the average heat current in the presence of non-Markovianity and briefly comment on the impact that alternative system-environment interaction terms can have on the steady-state properties.
@@ II Markovian Quantum Collision models
II.1 Single-bath thermalizing model (setting I)
@@
Collision models are a flexible framework for studying open system dynamics CICCARELLO20221; Campbell2021EPL; indeed they can be used to simulate any time-continuous open system dynamics described by a Markovian quantum master equation MEsfromCMs; Cattaneo2021collision. A simple setting of wide physical relevance is a two-level system with Hamiltonian, $H_{S}=\tfrac{\omega_{S}}{2}\sigma_{z}$, thermalising with a bosonic environment described by the master equation Current_fluctuations,
@@
where $\Gamma$ is the coupling strength and $\bar{N}\!=\!(e^{\beta\omega_{S}}-1)^{-1}$ is the Bose-Einstein distribution for a bath at (inverse) temperature $\beta\!=\!1/T$, as depicted in Fig. 1(a). The two dissipators $\mathcal{D}[\sigma^{-}],\mathcal{D}[\sigma^{+}]$ represent the emission and absorption of thermal photons, respectively, with $\sigma^{-}\!=\!|0\rangle\langle 1|$ and $\sigma^{+}\!=\!|1\rangle\langle 0|$ being the Pauli lowering and raising operators. The steady state of Eq. (4) is given by the Gibbs state,
@@
The dynamics described by Eq. (4) can be stroboscopically simulated using a collision model Rau1963; ScaraniPRL2002; CICCARELLO20221; cusumano2022quantum, where the environment is viewed as an infinite collection of auxiliary units, each modeled as a qubit, identically prepared in a thermal state, i.e. $\eta_{n}=\rho^{\text{th}}$ given by Eq. (5), with the free evolution of each unit given by $H_{A_{n}}=\tfrac{\omega_{A_{n}}}{2}\sigma^{z}$. For simplicity, in the rest of this work we shall always assume the system and auxiliary units are on resonance, i.e. $\omega_{S}\!=\!\omega_{A_{n}}\!\equiv\!\omega$. This setting is illustrated in Fig. 1(b). In order to match the dynamics given by Eq. (4), we fix the interaction Hamiltonian between the $S$ and $A_{n}$ to be,
$H_{I}=-\frac{J}{2}\left(\sigma^{x}_{S}\otimes\sigma^{x}_{A_{n}}+\sigma^{y}_{S}\otimes\sigma^{y}_{A_{n}}\right)\;,$ (6)
@@
with $J=\sqrt{{\frac{\Gamma(2\bar{N}+1)}{\Delta t}}}$. Scaling the interaction strength by $1/\sqrt{\Delta t}$ is necessary to ensure that Eq. (4) is recovered in the limit $\Delta t\to 0$. Throughout this work we will refer to this as ‘setting I’. Henceforth, we shall move to the interaction picture and neglect the free evolution of the system and auxiliary units. This can be done due to the fact that sum of the free Hamiltonians commute with the interaction Hamiltonian, i.e. $\left[H_{S}+H_{A_{n}},H_{I}\right]=0$  Composite_CM. This fact also means there is no switching work required to stoboscopically turn on and off the system-bath interaction DeChiaraNJP; Guarnieri2020PLA.
[…]
@@
This setup leads to thermalization, wherein the steady state of $S$ converges to the initial auxiliary state, i.e. $\rho_{\infty}\!=\!\eta_{i}$, for any choice of initial state for the auxiliary units $A_{n}$ Homogenisation. When $S$ sequentially interacts with $A_{n}$ via Eq. (6), its dynamics are described by a contractive completely positive map, guaranteeing convergence toward $\eta$ as the unique fixed point. This process is thermodynamically consistent and captures key features of thermalization. Importantly, in the limit of infinitely many collisions, the system not only reaches equilibrium but also asymptotically factorizes from the environment, resulting in a product state without residual system-environment correlations GiovannettiPRA2018.
While here we will focus on system-environment interactions of the form Eq. (6), in the Appendix we consider an alternative given by the isotropic Heisenberg Hamiltonian. Although this interaction also leads to homogenization in setting I, the collision model no longer reduces to the same thermalisation dynamics governed by the GKSL equation, Eq. (4cusumano2024structured. Indeed, while many of the main features discussed below are similarly exhibited for the Heisenberg interaction, we highlight a number of interesting differences that emerge when alternative methods for simulating the open dynamics are considered in both the Markovian and non-Markovian regimes.
@@ II.2 Two-bath model (setting II)
[…]
@@
[…]
@@ II.3 Steady States in the Markovian Limit
We start our analysis by noting that settings I and II both simulate the dynamics of a system in contact with a thermal bath. Indeed, at the level of the system only, the two settings exhibit the same dynamics and steady state that would be expected under the time continuous evolution given by Eq. (4). Furthermore, it is evident that they are equivalent in the zero temperature limit, $\beta\rightarrow\infty$. This is an obvious consequence of the fact that, in this limit, $J^{(1)}\!\to\!0$ and thus the two bath case reduces to the zero temperature limit of setting I. However, as alluded to previously, and as we will show explicitly below, setting II generally drives the system to a NESS.
Maintaining the system in a NESS requires heat currents to be flowing. Because the unitary is strictly energy preserving, there is no work associated with switching on and off the collisional unitaries DeChiaraNJP; Guarnieri2020PLA. Hence all energy changes can be considered as heat exchanges. Focusing on the zero temperature bath in setting II, within the collision model setting we can readily calculate this via
@@
where $\tilde{\rho}_{A_{n}^{(0)}}$ is the state of the $A_{n}^{(0)}$ auxiliary unit after the system-environment collision has taken place. A similar expression holds for the $A_{n}^{(1)}$’s and, in fact, we find is exactly equal and opposite to Eq. (9), which is a direct consequence of strict energy conservation. Fig. 2 shows the steady state heat flux into the environment, i.e. $\frac{\Delta\mathcal{Q}}{\Delta t}\!\equiv\!\frac{\Delta\mathcal{Q}^{(0)}(n)}{\Delta t}$ for $n\!\rightarrow\!\infty$. We see that it is non-zero for all values of $\Delta t$. Smaller values of $\beta$ require larger heat fluxes to maintain the NESS. This is in line with expectation since in this regime the system is being driven to a state with a high effective temperature. It is therefore far from equilibrium from both of the respective collisional baths leading to substantial energy exchanges. Larger $\beta$ leads to a reduced heat flux since this regime tends to the zero temperature single bath situation captured by setting I. It is worth noting that the heat flux vanishes for all choices of $\beta$ in setting I Guarnieri2020PLA.
Figure 2: Steady states in the Markovian limit. Non-equilibrium steady state heat current, $\Delta\mathcal{Q}/\Delta t$ for the initially low energy (cold) bath, $A_{n}^{(0)}$, with increasing collision time-step $\Delta t$ for different canonical temperatures $\beta\!=\!2$ (dashed, blue; lower temperature) and $\beta\!=\!0.5$ (solid, red; higher temperature). Other parameters: $\omega\!=\!1,~\Gamma\!=\!4$.